OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08.

OpenStudy (amistre64):

solve the given formula for "t"

OpenStudy (amistre64):

A = P(1 + rt) A/P = 1+ rt ; etc ....

OpenStudy (anonymous):

Ok, if i give you the answers can you help me understand it.

OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08. A. 1.5 B. 2 C. 2.4 D. 3

OpenStudy (amistre64):

your simply spose to replace the letters in the formula with the numbers that they are equal to and come up with a suitable solution

OpenStudy (amistre64):

If you have not learned how to manipulate formulas and equations yet, then you might want to revisit it since this material is clearly built upon those abilities.

OpenStudy (anonymous):

Well that's why im on here for help.

OpenStudy (amistre64):

I know, but in trying to answer a question asked, its useful for to know some of the mechanics underlying it. Otherwise an answer itself is meaningless.

OpenStudy (amistre64):

i typed up the inputed stuff in the other one, just need to run it thru google to pop out an answer i for it i think :)

OpenStudy (anonymous):

Ok well thank you for your help. Means a lot. Goodbye